OpenStudy (anonymous):

What is an expression for all possible angles coterminal with 30 degrees in general form?

OpenStudy (joannablackwelder):

30+360n, where n is an integer.

OpenStudy (anonymous):

How did you figure that out? =)

OpenStudy (joannablackwelder):

Coterminal angles are angles that land in the same "spot" on the unit circle.

OpenStudy (joannablackwelder):

So, when you go one full rotation (to get back to the same spot), you travel 360 degrees either clockwise or counterclockwise.

OpenStudy (joannablackwelder):

And you can spin as many times as you want.

OpenStudy (joannablackwelder):

That is why n can be any integer.

OpenStudy (joannablackwelder):

Does that explanation help at all?

OpenStudy (anonymous):

Yes. Thank you. =)

OpenStudy (joannablackwelder):

No worries

OpenStudy (anonymous):

How about for 3 pi / 2?

OpenStudy (joannablackwelder):

Same concept applies, but for radians 1 rotation is 2π. Any ideas, guesses?

OpenStudy (anonymous):

2 pi n + 3 pi/2?

OpenStudy (joannablackwelder):

Right, where n is an integer. Good job!

OpenStudy (anonymous):

Awesome!

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

Can I please ask you another question?

OpenStudy (joannablackwelder):

Sure.

OpenStudy (anonymous):

For each angle in standard position, determine one positive and negative angle that is coterminal with it. pi/3

OpenStudy (joannablackwelder):

If you can write a formula for any cotermial angle, then you would just need to plug in two integers for n to find 2 coterminal angles, one positive and one negative.

OpenStudy (anonymous):

also, the question with the 30 degree angle, does the expression lie in the domain\[-720 degrees \le \theta < 1080 \]?

OpenStudy (joannablackwelder):

That depends on what you choose for n.

OpenStudy (anonymous):

but is that expression applicable here in general?

OpenStudy (joannablackwelder):

Not unless you restrict your n to only give you values in that domain.

OpenStudy (anonymous):

then what can n be restricted to?

OpenStudy (joannablackwelder):

Whatever you need it to be. When n is 0, θ is 30, so that is in your domain. When n is -1, θ is -330, so that is good too.

OpenStudy (joannablackwelder):

When n is -2, θ is -690. Is that in the domain?

OpenStudy (anonymous):

It should be.

OpenStudy (anonymous):

For the other question and the coterminal angles.... pi/3, is there an easier way to solve it without plugging it in the formula?

OpenStudy (joannablackwelder):

Right. Could n be -3 and θ be in the domain?

OpenStudy (joannablackwelder):

Not that I know of off hand...

OpenStudy (anonymous):

yes, it can because theta can be greater than 720 degrees and my answer is -1,050

OpenStudy (joannablackwelder):

I thought it couldn 't be less than -720.

OpenStudy (anonymous):

\[-720 \le \theta < 1080\]

OpenStudy (joannablackwelder):

You don 't have to plug in every value though. You can skip if it seems you have enough room.

OpenStudy (anonymous):

It looks good though.

OpenStudy (anonymous):

For the other question and the coterminal angles.... pi/3, is there an easier way to solve it without plugging it in the formula?

OpenStudy (joannablackwelder):

You don 't have to plug in every value for n. You can skip if it seems you have enough room.